OpenStudy (anonymous):

HELP ME PLZ!!!!!!! If (6y-11)(6y+11)=ay^2-b, what is the value of a?

jhonyy9 (jhonyy9):

use formula of difference of two squares - do you know it ?

jhonyy9 (jhonyy9):

a^2 -b^2 = ?

OpenStudy (jdoe0001):

\(\bf (6y-11)(6y+11)\qquad \qquad recall\qquad (a-b)(a+b) = a^2-b^2 \\ \quad \\ (6y-11)(6y+11)\implies ?\)

OpenStudy (anonymous):

Is it 36?

OpenStudy (anonymous):

Or 6?

OpenStudy (jdoe0001):

you tell us

OpenStudy (jdoe0001):

what would \(\bf (6y-11)(6y+11)\qquad \qquad recall\qquad (a-b)(a+b) = a^2-b^2 \\ \quad \\ (6y-11)(6y+11)\implies ?\) be?

OpenStudy (abdullahm):

User got help here: http://openstudy.com/study#/updates/56da1b45e4b0df3d187bbb3e